OpenStudy (anonymous):

Let f(x) = x - 3 and g(x) = 2x + 4. What is f(g(x))

OpenStudy (anonymous):

@chaise That's correct. In general, f(g(x)) means that we take the law of g(x) and put it on the x's that the rule of f(x) has, i.e., we substitute every x in f(x) by g(x).

OpenStudy (chaise):

f(g(x)) f(2x+4) f(x)=x-3 f(2x+4) f(2x+4)=(2x+4)-3 f(2x+4)=2x+1 I hope this is right, been a while since I have dnoe this sort of stuff ^_^ Slight copying error, sorry sir.